If f(x) = 4x + 3, evaluate f(x+h).
If g(x) = x², evaluate g(x+h).
If f(x) = -x + 1, evaluate f(x+h).
If f(x) = -x + 1, evaluate f(x-h).
Use the quadratic formula to solve the equation: x² - x - 1 = 0.

Answer by stanbon(75887)

If f(x) = 4x + 3, evaluate f(x+h).
f(x+h) = 4(x+h)+3 = 4x + 4h + 3
-------------------------------------
If g(x) = x², evaluate g(x+h)
g(x+h) = (x+h)^2 = x^2 + 2hx + h^2
--------------------------------------
etc.
If f(x) = -x + 1, evaluate f(x+h).
If f(x) = -x + 1, evaluate f(x-h).
Use the quadratic formula to solve the equation:
x² - x - 1 = 0.
x = [1 +- sqrt(1-4*1*-1)]/(2) = (1+sqrt(5))/2 or (1-sqrt(5))/2
--------------
Cheers,
Stan H.
---------------
(